what is the linear equation in slope-itercept form for through (0,2) and (-2, 3)

Answer by friesr(113)

y = mx + b
-
m= (y2-y1)/(x2-x1)
-
m = 3 - 2 / -2 - 0
m = -1/2
--
y = -1/2x + b
Lets sub in -2 for x and 3 for y to solve for b
-
3 = -1/2(-2) + b
b = 2
-
y = -1/2x + 2
===